# Exploring Test-Time Adaptation for Object Detection in Continually Changing Environments

### v3

@@ 3.2.2. Object-Level Contrastive Learning
SimCLR is a widely used CL approach for self-supervised visual representation learning (Chen et al., 2020). SimCLR learns high-quality feature representations from differently augmented views of the same image. Designed for classification tasks, it requires large batch sizes to ensure sufficient positive and negative pairs. Consequently, SimCLR is not well-suited for object detection, where images contain multiple instances and large batch sizes demand significant computational resources.
Motivated by SimCLR (Chen et al., 2020), we present an OCL module to extract teacher and student features for CL based on proposals generated from the RPN. This strategy provides multiple cropped views around the object instance, eliminating the need for large batch sizes and ensuring computational efficiency for online detection model updates. Specifically, given a weakly augmented image $f_{weak}(x_{t})$ at time step t, the teacher produces ROI proposals $p_{t}=\{p_{1},...,p_{l}\}$ via region proposal network. We then apply RoIAlign (He et al., 2017), a pooling operation for ROI, to extract corresponding teacher and student object-level features $F_{t}^{T}=\{f_{i}^{T}\in\mathbb{R}^{1\times C}\}_{i=1}^{l}$ and $F_{t}^{S}=\{f_{i}^{S}\in\mathbb{R}^{1\times C}\}_{i=1}^{l}$ based on the feature map from the backbone, respectively. The features associated with the same proposal are considered positive pairs, otherwise negative pairs. we employ the CL loss to these features $F_{t}^{T}$ and $F_{t}^{S}$ derived from weak and strong enhanced images respectively by minimizing:
@@
where $\tau>0$ is the temperature, $l$ indicates the number of the features, and $f_{i}^{T}$ and $f_{i}^{S}$ denote the features of two different augmentations of the same object, serving as the positive pair. This strategy encourages the model to learn fine-grained and localized representations on the target domain, without relying on accurate pseudo-labels.
@@ 3.2.3. Adaptive Monitoring
The AM module monitors the model’s status using the predicted scores, featuring dynamic skipping and dynamic threshold. The former halts unnecessary adaptations to save computational resources, while the latter updates the category-specific threshold.
As shown in Figure 2, the teacher network first make predictions $\hat{y}_{t}=f_{\theta_{t-1}}(f_{wek}(x_{t}))$ for the weakly enhanced images. Subsequently, We compute the mean scores of all category $\overline{l}_{t}$ in teacher’s predictions and maintain its exponentially moving average $\overline{l}_{t}^{ema}$ through:
@@
For dynamic threshold, the thresholds for each category are initialized with the same value $\delta_{0}$ at first, which is updated every iteration by:
@@
3.2.4. Data-Driven Stochastic Restoration
Existing methods address catastrophic forgetting through stochastic reset mechanisms (Wang et al., 2022) or pure data-driven restoration (Brahma and Rai, 2023). Nevertheless, the former may reset valuable parameters, potentially erasing essential knowledge relevant to the current domain. On the other hand, the latter restores the most unimportant parameters, which may retain noise parameters, thereby leading to error accumulation. Therefore, we propose the DSR strategy to reset irrelevant parameters with higher possibilities, enabling robust adaptation. Following (Brahma and Rai, 2023), we adopt the diagonal approximation of Fisher Information Matrix (FIM) (Kirkpatrick et al., 2017) based on gradient, as a metric of parameter importance. The FIM approximates parameter importance by considering the sensitivity of the loss function to parameter changes (Kirkpatrick et al., 2017). Parameters with higher values in the FIM are those for which small changes result in significant increases in loss, indicating their importance for the domain at hand.
Specifically, consider a convolution layer with weight $W_{t-1}$ and its corresponding gradient matrix after gradient descent $G_{t}$ with the same shape as $W_{t-1}$ at time step $t$. First, we use a random matrix $R_{t}\sim\operatorname{Uniform}(0,1)$ with the same shape as the weights matrix, where each element follows a uniform distribution between 0 and 1. Moreover, we utilize the element-wise multiplication of the square of the gradient and the random matrix to approximate the FIM: $F_{t}=G_{t}\odot G_{t}\odot R_{t}$, where $\odot$ represents the element-wise multiplication. Furthermore, the proposed DSR updates the parameters based on the FIM values, which is formulated as:
@@ 3.2.4. Data-Driven Stochastic Restoration
(10) $W_{t}=M_{t}\odot W_{0}+(1-M_{t})\odot W_{t-1},$
where $M_{t}$ indicates the mask matrix of the same shape as $W_{t}$, $<$ represents the element-wise less than operation, and $\eta$ is the threshold value which is acquired by the $q$-quantile of $F_{t}$, i.e. $\eta=quantile(F_{t},q)$. Consequently, the elements in $M_{t}$ are set to 1, when the corresponding FIM value is less than $\eta$, indicating the corresponding parameters should be reset to the source weight $W_{0}$. This strategy enables the model to retain essential knowledge while introducing randomness to enhance robustness to mitigate forgetting.
@@ 4. EXPERIMENTS
Table 1. Experimental results (mAP@0.5) and adaptation iterations of Cityscapes-to-Cityscapes-C short-term CTTA task. We evaluate the performance by continually adapting the source model to the twelve corruptions. “-unstop” denotes no skipping.
@@
Table 2. Experimental results (mAP@0.5) and adaptation iterations of Cityscapes-to-Cityscapes-C long-term CTTA task. We evaluate the performance by continually adapting the source model to the five corruption ten times. To save space, we display selected rounds of results, and full results are provided in the appendix. “-unstop” denotes no skipping.
In this study, we rigorously evaluate our methodology across four benchmark tasks in CTTA object detection. These tasks encompass continual adaptation to synthetic and real-world distribution shifts, evaluated over short and long-term periods. Inspired by the foundational work CoTTA (Wang et al., 2022), the short-term CTTA task entails the sequential adaptation to various target domains once. In contrast, the long-term CTTA task involves continually adapting the model toward groups of target domains cyclically.
@@ 4.1. Datasets
Our short-term CTTA task selectively focuses on the latter 12 corruptions. For the long-term CTTA task, we prioritize the five corruptions related to autonomous driving scenarios as the target domain groups following (Gan et al., 2023), namely Fog, Motion, Snow, Brightness, and Defocus. We repeat the adaptation to the target domain group 10 times to evaluate long-term performance.
@@
ACDC. The ACDC dataset (Sakaridis et al., 2021) shares the same class types as Cityscapes and is collected in four different adverse visual conditions, including Fog, Night, Rain, and Snow. Following (Wang et al., 2022), we use these four conditions as the target domain group for the long-term CTTA task, with 400 unlabeled images per condition. Similarly, the source-trained model is continually adapted to the target domain group for 10 cycles.
@@ 4.2. Implementation Details
We adopt the Faster R-CNN with ResNet50 (He et al., 2016) pre-trained on ImageNet (Krizhevsky et al., 2012) as the backbone. Aligning with (VS et al., 2023b; Wang et al., 2022), we maintain a batch size of 1, thereby emulating a real-world application scenario where the detection model adapts toward a continuous influx of images. The source models are trained using an SGD optimizer with a learning rate of 0.001 and a momentum of 0.9. Algorithms are implemented leveraging the Detectron2 platform (Wu et al., 2019). The metric of mAP at an IoU threshold of 0.5 is employed for evaluation. Each experiment is conducted on 1 NVIDIA A800 GPU. More hyper-parameters and implementation details are provided in the appendix.
@@ 4.3. Baselines and Compared Approaches
We compare CTAOD with six baselines across various settings, including Source (Ren et al., 2015), IRG (VS et al., 2023a), Memclr (VS et al., 2023b), CoTTA (Wang et al., 2022), SVDP (Yang et al., 2024), and CTAOD-upstop. Specifically, ”Source” represents the source model without adaptation. Memclr and IRG are SFDA object detection methods, also referred to as TTA methods (Wang et al., 2022). MemCLR integrates cross-attention with CL, while IRG incorporates instance relation graph and CL. CoTTA utilizes weight-averaged predictions and random neurons recovery to tackle CTTA, while SVDP explores sparse visual prompts for CTTA dense prediction. “CTAOD-unstop” represents CTAOD without dynamic skipping.
@@ 4.4. Results and Analysis
Table 3. Experimental results (mAP@0.5) and adaptation iterations of Cityscapes-to-ACDC long-term CTTA task. We evaluate the performance by continually adapting the source model to the four conditions ten times. To save space, we display selected rounds of results, and full results are provided in the appendix. “-unstop” denotes no skipping.
@@
(c) CTAOD (ours)
Figure 3. Qualitative results. We compare the detection results of the Source model, MemCLR, and CTAOD in the $10^{th}$ round of adaption to defocus corruption on the long-term Cityscapes-to-Cityscapes-C task.
4.4.1. Synthetic Continual Distribution Shift
We first evaluate the effectiveness of CTAOD on the short-term Cityscapes-to-Cityscapes-C adaptation tasks. As depicted in Table 1, CoTTA, IRG, and Memclr exhibit enhancements in performance, achieving 16.0, 17.5, and 18.0 mAP respectively. Although the SFDA methods like IRG and MemCLR assume a static target domain, employing a large momentum update rate $\alpha$ for the teacher enables a relatively stable adaptation in the short term. Consequently, CL leads to a better performance than CoTTA which solely employs weight-averaged predictions. Moreover, SVDP, which utilizes visual prompts, achieves a sub-optimal performance of 19.1 mAP. Remarkably, our proposed method improves the performance to 20.8 mAP, consistently outperforming the above approaches. Additionally, the dynamic skipping strategy in AM reduces adaptation iterations by 20% while achieving a 0.2 mAP improvement compared to CTAOD-unstop. CTAOD ensures a more reliable and efficient adaptation to target domains characterized by intense and continual changes in the short term.
As presented in Table 8, the long-term task outcomes reveal the source model’s poor performance, with an average mAP of 16.4. Despite the improvement of Memclr and IRG, their performance also begins to decline in the later rounds, failing to maintain stability over long-term adaptation. We believe this is due to the aforementioned methods not accounting for continual distribution shifts, resulting in error accumulation and catastrophic forgetting. Furthermore, CoTTA utilizes a stochastic restoration mechanism to mitigate forgetting, but its randomness may result in losing crucial information, thus limiting its performance at 19.1 mAP. In contrast, SVDP which employs sparse visual prompts raises the performance to 25.3 mAP. Particularly, CTAOD yields a remarkable 12.8 mAP enhancement over the Source and surpasses all comparative baselines using 80% adaptation iterations, which employs the DSR mechanism to conserve valuable knowledge while eliminating noise from prior domains. These findings empirically validate the effectiveness of our proposed method in ensuring stable and efficient adaptation amidst synthetic continual distribution shifts, across both short-term and long-term scenarios.
@@ 4.4.2. Real-World Continual Distribution Shift
We also evaluate our method on the real-world continual distribution shift dataset. The results of the SHIFT short-term CTTA task are shown in Table 4. While other baseline methods suffer from performance decline, SVDP which employs a fixed threshold for pseudo-labeling and prompts learning achieves a 0.5 mAP improvement. Moreover, CTAOD which introduces a dynamic threshold strategy in AM shows a superior performance of 43.9 mAP with 16% improvement in efficiency.
The experimental results of ACDC long-term CTTA task are summarized in Table 3. CoTTA, IRG, and MemCLR potentially suffer from error accumulation and catastrophic forgetting, manifesting in a rapid decline in performance during the later stages. Similarly, SVDP achieves a sub-optimal performance of 39.5 mAP. Notably, CTAOD achieves a performance of 40.4 mAP, yielding an absolute improvement of 0.9 mAP over baselines. While CTAOD may not display a significant advantage over other methods in the initial round, its performance continually improves and remains stable throughout the long-term adaptation. These results underscore the capability of our method to foster robust adaptation toward the real-world continual distribution shift in both the short and long term.
4.4.3. Qualitative Results
As shown in Figure 3, we selectively compare the source model, MemCLR, and CTAOD in the final round adaptation to the brightness on the long-term Cityscapes-to-Cityscapes-C tasks. CTAOD guides the model in learning better feature representations while effectively mitigating forgetting in the long-term adaptation. Therefore, CTAOD assists the detector in distinguishing more foreground object categories and better locating them. More visualization analysis is presented in the appendix.
Mean-Teacher OCL AM DSR Mean Gain
@@ 4.4.3. Qualitative Results
Table 5. Alation experiment. “Mean-Teacher” represents the base mean teacher framework with weak-strong augmentation and KL divergence distillation. All experiments are done on long-term Cityscapes-to-Cityscapes-C tasks.
4.4.4. Ablation Study
We conduct an ablation study to empirically assess the impact of the OCL, AM, and DSR. As shown in Table 5, the integration of the mean-teacher architecture elevates the performance to 25.9 mAP. Building upon the mean-teacher foundation, the proposed OCL, AM, and DSR achieve an additional 1.5, 1.4, and 0.8 mAP improvement individually. Moreover, the combination of the OCL and AM modules raises the performance to 28.6 mAP. Furthermore, the inclusion of the DSR mechanism brings a 0.6 mAP boost in performance, culminating in a comprehensive model performance of 29.2 mAP. More fine-grained comparisons with AM and DSR are provided in the appendix.
5. CONCLUSION
In this work, we proposed CTAOD, aimed at addressing the challenges of low-quality pseudo-labels and random errors resulting from stochastic recovery in continual test-tiem adaptation. Firstly, object-level contrastive learning leverages teacher ROI features for contrastive learning to refine the feature representation, tailored for object detection. Secondly, the dynamic threshold strategy selects higher-quality pseudo-labels by dynamically updating the category-specific threshold based on the predicted confidence scores. Lastly, the data-driven stochastic restoration mechanism selectively reset inactive parameters to mitigate forgetting while retaining valuable knowledge. The empirical results of four continual test-time adaptation tasks for object detection demonstrate the efficacy of CTAOD in both short-term and long-term adaption scenarios.
Appendix A Appendix Overview
In this appendix, we provide additional details and analysis. In the first section, we provide implementation details on CTAOD and other baseline methodologies. In the second section C, we furnish complete results for the quantitative research and additional visualization examples for the qualitative research. In the third section, we further discuss the continual learning. In the last section E, we discuss the limitations and prospective directions of our research.
Appendix B Additional Implementation Details
Following (VS et al., 2023b), we trained a Faster R-CNN (Ren et al., 2015) with ResNet50 (He et al., 2016) pre-trained on ImageNet (Krizhevsky et al., 2012) on the training set of Cityscapes (Cordts et al., 2016) as the source-trained model of Cityscapes. We utilized an SGD optimizer with a learning rate of 0.001, a momentum of 0.9, and a batch size of 1 and 70k iterations. Similarly, we trained another source model on the clear condition of the SHIFT (Sun et al., 2022) with 30 epochs. We will specify the implementation details of the CTAOD and other baseline methods in the following subsections.
B.1. Implememtation of CTAOD
Detailed hyper-parameters of CTAOD for each benchmark are listed in Table 7. Furthermore, we present a pseudo-code outline of our CTAOD framework in Algorithm 1. The three key components of CTAOD are highlighted, i.e., Object-level Contrastive Learning, Adaptive Monitoring, and Data-driven Stochastic Restoration.
B.2. Implememtation of Baseline Methods
The CoTTA (Wang et al., 2022) and SVDP (Yang et al., 2024) are originally designed for the classification or segmentation tasks. We implement them based on their public code, with some modifications to fit into the objection detection. Specifically, we implement the Cotta method by constructing a mean-teacher framework based on Faster R-CNN, which incorporates a stochastic restoration mechanism. Similarly, we implement SVDP with visual prompts on the area of the bounding box in pseudo-labels based on mean-teacher framework. Further, the MemCLR (VS et al., 2023b) and IRG (VS et al., 2023a) are source-free domain adaptation methods for object detection. Therefore, we implement the MemCLR and IRG based on their public code, with little modifications to follow the CTTA setting (i.e., updating the model parameters during evaluation). For a fair comparison, we set the same momentum $\alpha$ of the Exponential Moving Average (EMA) for CoTTA, SVDP, MemCLR, and IRG as CTAOD, all of which apply the mean-teacher framework (Tarvainen and Valpola, 2017). Other hyper-parameters are retained as specified in their public code.
Appendix C Additional Experiment Results
C.1. Quantitative Results
We present the complete experiment results on the two long-term CTTA tasks, i.e., Cityscapes-to-Cityscapes-C and Cityscapes-to-ACDC adaptation tasks, in Table 8 and 9. Experiments show that the proposed CTAOD can largely maintain strong performance in the long term.
C.2. Qualitative Results
As shown in figure 4 - 8, we present more visualization results to compare the CTAOD and other baseline methods based on the final round adaption on the long-term Cityscapes-to-Cityscapes-C task. The results further validate the effectiveness of CTAOD in distinguishing more foreground object categories and better locating them, compared with other baseline methods.
C.3. Additional Ablation Study
CTAOD-unstop FT of 0.9 FT of 0.8 FT of 0.7 SR DR Mean

### v4

@@ 3.2.2. Object-Level Contrastive Learning
SimCLR is a widely used CL approach for self-supervised visual representation learning (Chen et al., 2020). It learns high-quality feature representations from differently augmented views of the same image, which requires large batch sizes to ensure sufficient positive and negative pairs for classification tasks. Consequently, SimCLR is less suitable for object detection, where images contain multiple instances and large batch sizes demand significant computational resources.
Motivated by SimCLR (Chen et al., 2020), we present an OCL module to extract teacher and student features for CL based on proposals from the RPN. This strategy provides multiple cropped views around the object instance, eliminating the need for large batch sizes and ensuring computational efficiency for online detector updates. Specifically, given a weakly augmented image $f_{weak}(x_{t})$ at time step t, the teacher produces ROI proposals $p_{t}=\{p_{1},...,p_{l}\}$ via region proposal network. OCL then apply RoIAlign (He et al., 2017) to extract corresponding teacher and student object-level features $F_{t}^{T}=\{f_{i}^{T}\in\mathbb{R}^{1\times C}\}_{i=1}^{l}$ and $F_{t}^{S}=\{f_{i}^{S}\in\mathbb{R}^{1\times C}\}_{i=1}^{l}$ based on the feature map from the backbone, respectively. The features associated with the same proposal are considered positive pairs, otherwise negative pairs. Then CL loss is applied to these features $F_{t}^{T}$ and $F_{t}^{S}$ derived from weak and strong enhanced images respectively by minimizing:
@@
where $\tau>0$ is the temperature, $l$ indicates the number of the features, and $f_{i}^{T}$ and $f_{i}^{S}$ denote the features of two different augmentations of the same object, serving as the positive pair. This strategy encourages the model to learn fine-grained and localized feature representations on the target domain, without relying on accurate pseudo-labels.
@@ 3.2.3. Adaptive Monitoring
The AM module monitors the model’s status using the predicted scores, featuring dynamic skipping and dynamic threshold. The former halts unnecessary adaptations to save computational resources, while the latter updates the category-specific threshold dynamically.
As shown in Figure 2, the teacher network first make predictions $\hat{y}_{t}=f_{\theta_{t-1}}(f_{wek}(x_{t}))$ for the weakly enhanced images. Subsequently, AM computes the overall mean scores across all predicted instances $\overline{l}_{t}$ in teacher’s predictions and maintains its exponentially moving average $\overline{l}_{t}^{ema}$ through:
@@
For dynamic threshold, the thresholds for each category are initialized with the same value $\delta_{0}$ at first, which will be updated every iteration by:
@@
3.2.4. Adaptive Randomized Restoration
Existing methods address catastrophic forgetting through stochastic reset (Wang et al., 2022) or pure data-driven restoration (Brahma and Rai, 2023). Nevertheless, the former may reset valuable parameters, potentially erasing essential knowledge relevant to the current domain. On the other hand, the latter restores the most unimportant parameters, which may retain noise parameters, thereby leading to error accumulation. Unlike PETAL (Brahma and Rai, 2023), which was designed for classification tasks and resets the least active parameters across all layers, our approach refines this for object detection by identifying inactive parameters within each layer individually. This layer-specific technique accounts for the varying parameter scales at different depths of the model, and by further incorporating randomness, we propose ARR to reset irrelevant parameters with higher possibilities to enhance stability and robustness in restoration. Following (Brahma and Rai, 2023), we adopt the diagonal approximation of Fisher Information Matrix (FIM) (Kirkpatrick et al., 2017) based on gradient, as a metric of parameter importance. The FIM approximates parameter importance by considering the sensitivity of the loss function to parameter changes (Kirkpatrick et al., 2017). Parameters with higher values in the FIM are those for which small changes result in significant increases in loss, indicating their importance for the domain at hand.
Specifically, given a layer with weight $W_{t}$ updated through gradient descent to $\hat{W}_{t}$ based on gradient matrix $G_{t}$ at time step $t$, ARR generates a random matrix $R_{t}\sim\operatorname{Uniform}(0,1)$ with the same shape as the weights matrix, where each element follows a uniform distribution between 0 and 1. Then the element-wise multiplication of the squared gradient matrix and the random matrix is applied to approximate the FIM. Furthermore, ARR updates the parameters based on the FIM values, formulated as:
(9) $F_{t}=G_{t}\odot G_{t}\odot R_{t},$
@@ 3.2.4. Adaptive Randomized Restoration
(11) $W_{t+1}=M_{t}\odot W_{0}+(1-M_{t})\odot\hat{W}_{t},$
where $\odot$ is the element-wise multiplication, $M_{t}$ indicates the mask matrix, $<$ represents element-wise less than operation, and $\eta$ is the threshold value acquired by the $q$-quantile of $F_{t}$: $\eta=quantile(F_{t},q)$. Consequently, the elements in $M_{t}$ are set to 1, when the corresponding FIM value is less than $\eta$, indicating the corresponding parameters should be reset to the source weight $W_{0}$. This strategy enables the model to retain essential knowledge while introducing randomness to enhance robustness to mitigate forgetting.
@@ 4. EXPERIMENTS
Table 1. Experimental results (mAP@0.5) and adaptation iterations of Cityscapes-to-Cityscapes-C short-term CTTA task. We evaluate the performance by continually adapting the source model to the twelve corruptions. “-unstop” denotes AMROD without skipping.
@@
Table 2. Experimental results (mAP@0.5) and adaptation iterations of Cityscapes-to-Cityscapes-C long-term CTTA task. We evaluate the performance by continually adapting the source model to the five corruption ten times. “-unstop” denotes AMROD without skipping. To save space, we display selected rounds of results, and full results are provided in the supplementary material.
In this study, we rigorously evaluate our methodology across four benchmark tasks in CTTA object detection. These tasks encompass continual adaptation to synthetic and real-world distribution shifts, evaluated over short and long-term periods. Inspired by the foundational work CoTTA (Wang et al., 2022), the short-term CTTA task entails the sequential adaptation to various target domains once. In contrast, the long-term CTTA task involves continually adapting the model toward a group of target domains cyclically.
@@ 4.1. Datasets
Our short-term CTTA task selectively focuses on the latter 12 corruptions. For the long-term CTTA task, we prioritize the five corruptions related to autonomous driving scenarios as the target domain group following (Gan et al., 2023b), namely Fog, Motion, Snow, Brightness, and Defocus. We repeat the adaptation to the target domain group 10 times to evaluate long-term performance.
@@
ACDC. The ACDC (Sakaridis et al., 2021) shares the same class types as Cityscapes and is collected in four different adverse visual conditions, including Fog, Night, Rain, and Snow. Following (Wang et al., 2022), we use these four conditions as the target domain group for the long-term CTTA task, with 400 unlabeled images per condition. Similarly, the source model is continually adapted to the target domain group for 10 cycles.
@@ 4.2. Implementation Details
We adopt the Faster R-CNN with ResNet50 (He et al., 2016) pre-trained on ImageNet (Krizhevsky et al., 2012) as the backbone. Aligning with (VS et al., 2023b; Wang et al., 2022), we maintain a batch size of 1 to emulate a real-world application scenario where the detection model adapts toward a continuous influx of images. The source models are trained using an SGD optimizer with a learning rate of 0.001 and a momentum of 0.9. Algorithms are implemented leveraging the Detectron2 platform (Wu et al., 2019). The metric of mAP at an IoU threshold of 0.5 (mAP@0.5) is employed for evaluation. Each experiment is conducted on 1 NVIDIA A800 GPU. More hyper-parameters and implementation details are provided in the supplementary material.
@@ 4.3. Baselines and Compared Approaches
We compare AMROD with six baselines across various settings, including Source (Ren et al., 2015), IRG (VS et al., 2023a), Memclr (VS et al., 2023b), CoTTA (Wang et al., 2022), SVDP (Yang et al., 2024), and AMROD-upstop. Specifically, ”Source” represents the source model without adaptation. Memclr and IRG are SFDA object detection methods, also referred to as TTA methods (Wang et al., 2022). MemCLR integrates cross-attention with CL, while IRG incorporates instance relation graph and CL. CoTTA utilizes a fixed threshold for pseudo-labeling and random neuron recovery to tackle CTTA, while SVDP explores sparse visual prompts for CTTA dense prediction. “AMROD-unstop” represents AMROD without dynamic skipping.
@@ 4.4. Results and Analysis
4.4.1. Synthetic Continual Distribution Shift
We first evaluate the effectiveness of AMROD on the short-term Cityscapes-to-Cityscapes-C adaptation tasks. As depicted in Table 1, CoTTA, IRG, and Memclr exhibit enhancements in performance, achieving 16.0, 17.5, and 18.0 mAP respectively. Although the SFDA methods like IRG and MemCLR assume a static target domain, employing a large momentum update rate $\alpha$ for the teacher enables a relatively stable adaptation in the short term. Consequently, CL leads to a better performance than CoTTA which solely employs weight-averaged predictions. Moreover, SVDP, which utilizes visual prompts, achieves a sub-optimal performance of 19.1 mAP. Remarkably, our proposed method improves the performance to 20.8 mAP, consistently outperforming the above approaches. Additionally, the dynamic skipping strategy in AM reduces adaptation iterations by 20% while achieving a 0.2 mAP improvement compared to AMROD-unstop. AMROD ensures a more reliable and efficient adaptation to target domains characterized by intense and continual changes in the short term.
@@ 4.4.1. Synthetic Continual Distribution Shift
Table 3. Experimental results (mAP@0.5) and adaptation iterations of Cityscapes-to-ACDC long-term CTTA task. We evaluate the performance by continually adapting the source model to the four conditions ten times. “AMROD-unstop” represents AMROD without dynamic skipping. To save space, we display selected rounds of results, and full results are provided in the supplementary material.
@@
(c) AMROD (ours)
Figure 3. Qualitative results. We compare the detection results of the Source model, MemCLR, and AMROD in the $10^{th}$ round of adaption to defocus corruption on the long-term Cityscapes-to-Cityscapes-C task.
As presented in Table 8, the long-term task outcomes reveal the source model’s poor performance, with an average mAP of 16.4. Despite the improvement of Memclr and IRG, their performance also begins to decline in the later rounds, failing to maintain stability over long-term adaptation. We believe this is due to the aforementioned methods not accounting for continual distribution shifts, resulting in error accumulation and catastrophic forgetting. Furthermore, CoTTA utilizes a stochastic restoration mechanism to mitigate forgetting, but its randomness may result in losing crucial information, thus limiting its performance at 19.1 mAP. In contrast, SVDP which employs sparse visual prompts raises the performance to 25.3 mAP. Particularly, AMROD yields a remarkable 12.8 mAP enhancement over the Source and surpasses all comparative baselines using 80% adaptation iterations, which employs the ARR mechanism to conserve valuable knowledge while eliminating noise from prior domains. These findings empirically validate the effectiveness of our proposed method in ensuring stable and efficient adaptation amidst synthetic continual distribution shifts, across both short-term and long-term scenarios.
@@ 4.4.2. Real-World Continual Distribution Shift
We also evaluate our method on the real-world continual distribution shift dataset. The results of the SHIFT short-term CTTA task are shown in Table 4. While other baseline methods suffer from performance decline, SVDP which employs a fixed threshold for pseudo-labeling, stochastic restoration, and prompts learning achieves a 0.5 mAP improvement. Moreover, AMROD which introduces AM and ARR modules shows a superior performance of 43.9 mAP with 16% improvement in efficiency.
The experimental results of the ACDC long-term CTTA task are summarized in Table 3. CoTTA, IRG, and MemCLR potentially suffer from error accumulation and catastrophic forgetting, manifesting in a rapid decline in performance during the later stages. Similarly, SVDP achieves a sub-optimal performance of 39.5 mAP. Notably, AMROD achieves a performance of 40.4 mAP, yielding an absolute improvement of 0.9 mAP over baselines. While AMROD may not display a significant advantage over other methods in the initial round, its performance continually improves and remains stable throughout the long-term adaptation. These results underscore the capability of our method to foster robust adaptation toward the real-world continual distribution shift in both the short and long term.
Mean-Teacher OCL AM ARR Mean Gain
@@
Table 5. Main ablation experiment. “Mean-Teacher” represents the base mean teacher framework with weak-strong augmentation and KL divergence distillation. All experiments are done on long-term Cityscapes-to-Cityscapes-C tasks.
4.4.3. Qualitative Results
As shown in Figure 3, we selectively compare the source model, MemCLR, and AMROD in the final round adaptation to the brightness on the long-term Cityscapes-to-Cityscapes-C tasks. AMROD guides the model in learning better feature representations while effectively mitigating forgetting in the long-term adaptation. Therefore, AMROD assists the detector in distinguishing more foreground object categories and better locating them. More visualization analysis is presented in the supplementary material.
AMROD FT of 0.9 FT of 0.8 FT of 0.7 SR DR Mean